OpenStudy (anonymous):

Identify function f(x) and g(x) such that the given function equals f[g(x)]? a) (1/x^2)+1 help me please..

OpenStudy (anonymous):

g(x) = 1/x^2 f(x) = x +1

OpenStudy (anonymous):

how u solve it?teach me please

OpenStudy (anonymous):

ah actually you can do it in more than one way, I mean we don't have single solution here for example f(x) = x g(x) = (1/x^2) +1

OpenStudy (anonymous):

meaning that it may have more than one solution?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

if f(x) = x and g(x) = x^2 +1 f[g(x)] = x^2 +1 now all you have to do is inverse of this process

OpenStudy (anonymous):

to get the f(x) and g(x)

OpenStudy (anonymous):

ouhh..so its depend on how we solve it?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

Here g is inside f. g is a function of x (namely 1/x^2). f is another function of something (namely something + 1).